Please help!
$1500 deposited at the beginning of each year for 15 years at 5% compounded annually. Find the future value of the annuity due.

Answer by rfer(12647)

FV=(P((1+i)^n-1))/(i)(1-i)
FV=(1500((1.05)^15-1)/(0.05)(0.95)
FV=(1500((2.079)-1)/(0.05)(0.95)
FV=1500*1.079)/(0.05)(0.95)
FV=1618.50/0.05(0.95)
FV=32370(0.95)
FV=$30,751.50